A gardener has six rows available in his vegetable garden to place tomatoes, eggplants, peppers, cucumbers, beans, and lettuce. Each vegetable will be allowed one and only one row. How many ways are there to position these vegetables in the garden?
Answer by ikleyn(52780)

.
A gardener has six rows available in his vegetable garden to place tomatoes, eggplants, peppers, cucumbers, beans, and lettuce.
Each vegetable will be allowed one and only one row. How many ways are there to position these vegetables in the garden?
~~~~~~~~~~~~~~~~


For any 6 different items, there are 6! = 6*5*4*3*2*1 = 720 different permutations.

For 6 different vegetables, there are 6! = 6*5*4*3*2*1 = 720 different ways to position them in the 6 rows of the garden.

--------------

Solved and explained.


This problem is on PERMUTATIONS.

On  Permutations,  see introductory lessons
    - Introduction to Permutations
    - PROOF of the formula on the number of Permutations
    - Simple and simplest problems on permutations
    - Special type permutations problems
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic  "Combinatorics: Combinations and permutations".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.